# A Survey on LLM-as-a-Judge

### v4

@@ 2.3. Post-processing Method
Post-processing refines the probability distributions generated by LLM-as-a-Judge to provide accurate evaluations. The evaluation format should align with our In-Context Learning design. Additionally, post-processing may involve procedures to enhance the reliability of extracted evaluations, closely linked to the In-Context Learning framework and consistently applied. There are three main methods of post-processing, which are extracting specific tokens, normalizing the output logits, and selecting sentences with high returns.
We will provide a detailed explanation of these methods. However, it is important to note that each method has significant limitations when evaluating objective questions. For example, in text response evaluation (Yu et al., 2024b), failing to accurately extract the key answer token from the LLM’s response can result in incorrect evaluation outcomes. These challenges in post-processing are deeply connected to the prompt design used in the earlier ICL stages and the selected model’s ability to follow instructions reliably.
@@ 2.3.3. Normalizing the output logits
In addition, Self-evaluation (Hao et al., 2023a) is also common using this method for LLM-as-a-Judge. It can be helpful to let the LLM evaluate itself by asking, ”Is this reasoning step correct?” and then reward it based on the probability of the next word being ”Yes.”
@@ 2.3.4. Selecting sentences
In addition to selecting specific tokens and normalizing the output logits, the content extracted by LLM-as-a-Judge may also be a sentence or paragraph. As showed in Figure 2, agent for reasoning task (Hao et al., 2023a), builds a reasoning tree by iteratively considering the most promising reasoning steps (actions, sub-questions) by LLM-as-a-Judge.
@@ 2.4. Evaluation Pipeline
After completing the three processes, we obtain the final evaluation $\mathcal{E}$. From input to output, these steps collectively form the LLM-as-a-Judge evaluation pipeline, as illustrated in Figure 2. This pipeline is commonly applied in four scenarios showed in Figure 14: LLM-as-a-Judge for models, LLM-as-a-Judge for data, LLM-as-a-Judge for agents, and LLM-as-a-Judge for reasoning/thinking.
@@ 2.4.1. LLM-as-a-Judge for Models
It is universally known that the best way to evaluate LLMs is human judgment, but collecting human annotations can be costly, time-consuming, and laborious (Ouyang et al., 2022b; Zheng et al., 2023d). Using strong LLMs (usually closed-source ones, e.g., GPT-4, Claude, ChatGPT) as an automated proxy for assessing LLMs has become a natural choice (Zhou et al., 2023a). With appropriate prompt design, the quality of evaluation and agreement to human judgment can be promising (Dubois et al., 2023; Zheng et al., 2023d; Zhang et al., 2023c; Wang et al., 2023b). However, the cost concern still exists when calling the APIs of these proprietary models, especially when there is a frequent need for model validation on large-scale data. Moreover, closed-source LLM-as-a-Judge leads to low reproducibility due to potential changes in models behind the API. Some recent works have started to make attempts for open-source alternatives. SelFee (Ye et al., 2023) collects generations, feedback, and revised generations from ChatGPT and fine-tunes LLaMA models to build a critique model. Shepherd (Wang et al., 2023c) trains a model that can output critiques for single-response with the data of feedback from online communities and human annotation. PandaLM (Wang et al., 2023d) trains a model to conduct pairwise comparison for LLM Instruction Tuning Optimization, and Zheng et al. (2023d) also fine-tune Vicuna (Touvron et al., 2023b) on a 20K pairwise comparison dataset to explore the potential of open-source models as a more cost-friendly proxy.
Figure 9. The illustrations of the scenario LLM-as-a-Judge for Models
@@ 2.4.2. LLM-as-a-Judge for Data
Data annotation generally refers to the labeling or generating of raw data with relevant information, which could be used for improving the efficacy of machine learning models. The process, however, is labor-intensive and costly. The emergence of LLMs presents an unprecedented opportunity to automate the complicated process of data annotation by LLM-as-a-Judge. Most of the data need to be evaluated by LLM-as-a-Judge is generated by models, or large-scale crawled data. Language models first conduct supervised fine-tuning to imitate how to align with human instructions (Wang et al., 2023a; Taori et al., 2023). After that, reinforcement learning techniques have been explored to align language models with human preferences (Ouyang et al., 2022c; Ramamurthy et al., 2023). The most successful way is applying a RLHF framework (Ouyang et al., 2022c) via training a reward model on human feedback and using PPO (Schulman et al., 2017) to obtain the policy model for language generation. However, in practices, the PPO training paradigm is complex in coding and hyper-parameter tuning while it needs four models that are hard for training. This motivates us to explore simpler and more straightforward methods to align language models with human preferences. This involves how to use LLM-as-a-Judge to evaluate whether different responses are aligned with human preferences. For example,  (Yuan et al., 2023; Dong et al., 2023) use general LLM (ChatGPT) to get better alignment with human preferences. The Aplaca prompts (Taori et al., 2023) is used as sampling queries to different models generate responses. And these data was evaluated by LLM-as-a-Judge to obtain human preference scores (reward score) to train a new language model. Other works would like to use Supervised Fine-Tuning (SFT) model itself as evaluator, like generating better-aligned datasets for SFT including hindsight-modified prompts (Zhang et al., 2023a; Liu et al., 2023c) and principle-driven self-alignment (Sun et al., 2023c).
In addition, the lack of domain-specific model training data is a common phenomenon. In order to obtain annotated high-quality data, it is also very common to use LLM-as-a-Judge for the generation and evaluation of domain data. WizardMath (Luo et al., 2023a) would use its Instruction Reward Model (IRM) as Evaluator, aiming to judge the quality of the evolved instructions on three aspects: i) Definition, ii) Precision, and iti) Integrity. To produce the ranking list training data of IRM, for each instruction, ChatGPT and Wizard-E are used to generate 2-4 evolved instructions respectively. Then we leverage Wizard-E to rank the quality of those 4-8 instructions.
@@
Recent research on evaluating multimodal data focuses on addressing vision-language misalignments in Multimodal Large Language Models (MLLMs), which often cause hallucinations—outputs inconsistent with visual or contextual evidence (Li et al., 2023a; Wang et al., 2023e; Cui et al., 2023). Techniques like Reinforcement Learning from Human Feedback (RLHF) and Factually Augmented RLHF have been employed to improve model alignment by incorporating structured ground-truth data and image captions, enhancing hallucination detection (Sun et al., 2023b). Benchmarks such as MLLM-as-a-Judge (Chen et al., 2024b) assess these models using tasks like scoring, pair comparison, and batch ranking, revealing limitations in alignment with human preferences. Persistent issues include biases (e.g., position, verbosity) and hallucinations, with even advanced models like GPT-4V displaying challenges. While pair comparison tasks align better with human judgment, scoring and batch ranking require significant improvements for reliable deployment. These findings emphasize the need for innovative frameworks and datasets to refine MLLM evaluation and alignment.
@@ 2.4.3. LLM-as-a-Judge for Agents
There are two ways to apply LLM-as-a-Judge for an agent. One is to evaluate the entire process of the intelligent agent (Zhuge et al., 2024), and the other is to evaluate it at a specific stage in the agent framework process (Hao et al., 2023a; Shinn et al., 2023). Both approaches are briefly illustrated in Figure 10. Using LLM as the brain of agent, an agentic system (Zhuge et al., 2024) could evaluate like a human, it would reduce the need for human involvement and eliminate the trade-off between thoroughness and effort. In addition, the agent (Shinn et al., 2023) can interact with the environment through language and receive feedback on actions through LLM to make decisions for the next action.
@@ 2.4.4. LLM-as-a-Judge for Reasoning/Thinking
The role of LLM-as-a-Judge in enhancing reasoning or thinking can be understood through two frameworks: scaling training time (Gao et al., 2023d; Trung et al., 2024) and scaling test time (Snell et al., 2024). In the training phase, LLM-as-a-Judge frequently operates within reinforcement learning paradigms, where it functions as a reward model or evaluator for data or processes. This enables the creation of high-quality reasoning datasets through mechanisms such as step-by-step verification (Lightman et al., 2023a), Direct Preference Optimization(DPO) (Rafailov et al., 2024), and self-refinement (Yuan et al., 2024). Recently, several LLMs trained with reinforcement learning to exhibit advanced reasoning and thinking abilities have gained attention, such as o1222https://openai.com/index/learning-to-reason-with-llms/, DeepSeek-R1333https://api-docs.deepseek.com/news/news1120,gemini-thinking [^4] https://ai.google.dev/gemini-api/docs/thinking-mode, and QVQ [^5] https://huggingface.co/Qwen/QVQ-72B-Preview. In the test-time framework, LLM-as-a-Judge is crucial for evaluating and selecting the best reasoning paths. For example, in ”Best-of-N” generation scenarios, where multiple reasoning outputs are produced, the judge determines the most accurate and coherent response. This dual role in both training and test phases demonstrates the indispensable nature of LLM-as-a-Judge in enhancing reasoning systems.
@@ 3. Improvement Strategy
When directly utilizing LLMs to conduct evaluation tasks such as scoring, selection, pairwise comparison, or ranking, the inherent biases of LLMs like length bias, position bias, and concreteness bias(Park et al., 2024) will lead to poor evaluation results. Addressing these inherent biases and improving the overall evaluation performance of LLMs is a critical challenge for applying LLMs as evaluators. In this section, we introduce three improvement strategies aimed at enhancing the evaluation performance of LLM-as-a-judge: design strategy of evaluation prompts (in-context learning based), improvement strategy of LLMs’ evaluation capabilities (model-based), and optimization strategy of final evaluation results (post-processing based). Our categorization is based on the formal definition of LLM-as-a-judge in Section 2, focusing on enhancing the evaluation effectiveness by targeting three key phases of the process: the context $\mathcal{C}$, the abilities of LLMs themselves $\mathcal{P_{LLM}}$ and the post-processing $\leftarrow$ to obtain the final results $\mathcal{E}$
@@
Evaluation Tasks, text width=6em [Few-shot prompting: FActScore (Min et al., 2023) / SALAD-Bench (Li et al., 2024c) / GPTScore (Fu et al., 2024)] [Evaluation steps decomposition: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024f) / SocREval (He et al., 2024) / BSM (Saha et al., 2024)] [Evaluation criteria decomposition: HD-Eval (Liu et al., 2024a) / Hu and Gao et al. (Hu et al., 2024)] [Shuffling contents: Wang et al. (Wang et al., 2024c) / Auto-J (Li et al., 2023b) / JudgeLM (Zhu et al., 2023a) / PandaLM (Wang et al., 2023d)] [Conversion of evaluation tasks: Liu et al. (Liu et al., 2024b)] ] [Optimizing LLMs’
Output Forms, text width=6em [Constraining outputs in structured formats: G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024f) / LLM-EVAL (Lin and Chen, 2023)] [Providing evaluations with explanations: CLAIR (Chan et al., 2023b) / FLEUR (Lee et al., 2024b)] ] ] [Improvement Strategy
@@
Evaluation Datasets, text width=7em [PandaLM (Wang et al., 2023d) / SALAD-Bench (Li et al., 2024c) / OffsetBias (Park et al., 2024) / JudgeLM (Zhu et al., 2023a) / CritiqueLLM (Ke et al., 2024)] ] [Iterative Optimization
Based on Feedbacks, text width=7em [INSTRUCTSCORE (Xu et al., 2023) / JADE (Zhang et al., 2023b)] ] ] [Optimization Strategy
@@
Evaluation Results, text width=7.5em [Summarize by multiple rounds: Sottana et al. (Sottana et al., 2023) / PsychoBench (Huang et al., 2023) / Auto-J (Li et al., 2023b)] [Vote by multiple LLMs: CPAD (Liu et al., 2023d) / Bai et al. (Bai et al., 2023b)] ] [Direct Optimization of
LLMs’ Outputs, text width=7.5em [Score smoothing: FLEUR (Lee et al., 2024b) / G-Eval (Liu et al., 2023a) / DHP (Wang et al., 2024f)] [Self validation: TrueTeacher (Gekhman et al., 2023)] ] ] ]
@@ 3.1. Design Strategy of Evaluation Prompts
An evaluation prompt is an input to LLM evaluators, which is used to guide the LLMs to complete the required evaluation tasks. LLMs possess in-context learning ability to learn how to perform specified tasks through relevant examples or instructions provided in prompts without requiring weight updates or retraining(Brown et al., 2020). It indicates that the design strategy of evaluation prompts will significantly impact the effectiveness of LLM-as-a-judge. Therefore, how to optimize the design of evaluation prompts, including better methods to help LLMs understand the evaluation tasks and produce evaluation results, is the most direct and effective way to improve the evaluation performance of LLM-as-a-judge.
@@ 3.1.1. Optimizing LLMs’ Understanding of Evaluation Tasks
In optimization methods of prompting LLMs to better understand evaluation tasks, one of the most commonly used and effective approaches is few-shot prompting(Brown et al., 2020). By incorporating several high-quality evaluation examples into the evaluation prompts, LLM evaluators can effectively grasp the objectives, general processes, and rough evaluation criteria of evaluation tasks. Many research works employ this prompt paradigm for evaluation, such as FActScore(Min et al., 2023), SALAD-Bench(Li et al., 2024c), and GPTScore(Fu et al., 2024).
@@
In addition to providing high-quality examples for LLMs, refining the evaluation task instructions is also an effective approach to optimize LLMs’ understanding of evaluation tasks. Current methods for refining evaluation tasks mainly include the decomposition of evaluation steps and criteria: (a) Decomposition of Evaluation Steps entails breaking down the entire evaluation tasks into smaller steps, providing detailed definitions and constraints for each small step in prompts, thereby guiding LLMs comprehensively through the whole evaluation pipeline. For instance, G-Eval(Liu et al., 2023a) and DHP(Wang et al., 2024f) use Chain-of-Thought(CoT)(Wei et al., 2022a) to guide LLMs. SocREval(He et al., 2024) employs the Socratic method to meticulously design each step to enhance evaluation performance. Saha et al. propose Branch-Solve-Merge(BSM)(Saha et al., 2024), which divides evaluation tasks into multiple parallel sub-tasks for separate evaluation and final merge. (b) Decomposition of Evaluation Criteria involves breaking down coarse evaluation criteria like Fluency into finer-grained sub-criteria like Grammar, Engagingness, and Readability, and then generating overall scores based on these difference dimensions. HD-Eval(Liu et al., 2024a) iteratively aligns LLM evaluators with human preference via hierarchical criteria decomposition thereby addressing the potential bias in LLMs. Hu and Gao et al.(Hu et al., 2024) summarize and clearly define an explicit hierarchical classification system encompassing 11 criteria, addressing the issue of LLMs potentially confusing different evaluation standards. These refinements are specific to enable LLMs to understand the details of evaluation tasks more deeply, thereby aligning evaluation results more closely with human evaluation requirements and preferences.
@@
Furthermore, the evaluation capabilities can be optimized based on specific shortcomings of LLMs in prompts. For instance, to address specific biases like position bias which is common in pairwise evaluations, several research efforts have optimized prompts design by randomly swapping contents to be evaluated. Wang et al.(Wang et al., 2024c) analyzed and validated the impact of position bias on LLM-as-a-judge and proposed a calibration framework to mitigate this bias by swapping the contents and averaging the scores. Auto-J(Li et al., 2023b) and JudgeLM(Zhu et al., 2023a) also enhance the evaluation consistency by shuffling the texts to be evaluated. In contrast to averaging scores, PandaLM(Wang et al., 2023d) annotates the conflicting evaluation results after swapping as ”Tie” to address the position bias.

### v5

@@ 2.3. Post-processing Method
Post-processing refines the probability distributions generated by LLM-as-a-Judge to ensure accurate evaluations. The evaluation format should align with our In-Context Learning design and may involve procedures to enhance the reliability of extracted evaluations, which should be applied consistently. We focus on three main post-processing methods: extracting specific tokens, normalizing the output logits, and selecting sentences with high returns.
However, it is important to note that each method has significant limitations when evaluating objective questions. For example, in text response evaluation (Yu et al., 2024b), failing to accurately extract the key answer token from the LLM’s response can result in incorrect evaluation outcomes. These challenges in post-processing are tightly linked to the prompt design used in earlier ICL stages and the selected model’s ability to follow instructions reliably.
@@ 2.3.3. Normalizing the output logits
In addition, Self-evaluation (Hao et al., 2023) is also common using this method for LLM-as-a-Judge. It can be helpful to let the LLM evaluate itself by asking, ”Is this reasoning step correct?” and then reward it based on the probability of the next word being ”Yes.”
@@ 2.3.4. Selecting sentences
In addition to selecting specific tokens and normalizing the output logits, the content extracted by LLM-as-a-Judge may also be a sentence or paragraph. As showed in Figure 2, agent for reasoning task (Hao et al., 2023), builds a reasoning tree by iteratively considering the most promising reasoning steps (actions, sub-questions) by LLM-as-a-Judge.
@@ 2.4. Evaluation Pipeline
After completing the three processes, we obtain the final evaluation $\mathcal{E}$. From input to output, these steps collectively constitute the LLM-as-a-Judge evaluation pipeline, as illustrated in Figure 2. This pipeline is commonly applied in four scenarios shown in Figure 8: LLM-as-a-Judge for models, LLM-as-a-Judge for data, LLM-as-a-Judge for agents, and LLM-as-a-Judge for reasoning or thinking.
@@ 2.4.1. LLM-as-a-Judge for Models
Figure 9. The illustrations of the scenario LLM-as-a-Judge for Models. The example of ”win-tie-lose” is from Li et al. (2023b)
It is universally known that the best way to evaluate LLMs is human judgment, but collecting human annotations can be costly, time-consuming, and laborious (Ouyang et al., 2022; Zheng et al., 2023b). Using strong LLMs (usually closed-source ones, e.g., GPT-4, Claude, ChatGPT) as an automated proxy for assessing LLMs has become a natural choice (Zhou et al., 2023a), as shown in Figure 9. With appropriate prompt design, the quality of evaluation and agreement to human judgment can be promising (Dubois et al., 2023; Zheng et al., 2023b; Zhang et al., 2023d; Wang et al., 2024c). However, the cost concern still exists when calling the APIs of these proprietary models, especially when there is a frequent need for model validation on large-scale data. Moreover, closed-source LLM-as-a-Judge leads to low reproducibility due to potential changes in models behind the API. Some recent works have started to make attempts for open-source alternatives. SelFee (Ye et al., 2023) collects generations, feedback, and revised generations from ChatGPT and fine-tunes LLaMA models to build a critique model. Shepherd (Wang et al., 2023b) trains a model that can output critiques for single-response with the data of feedback from online communities and human annotation. PandaLM (Wang et al., 2023c) trains a model to conduct pairwise comparison for LLM Instruction Tuning Optimization, and Zheng et al. (2023b) also fine-tune Vicuna (Touvron et al., 2023b) on a 20K pairwise comparison dataset to explore the potential of open-source models as a more cost-friendly proxy.
@@ 2.4.2. LLM-as-a-Judge for Data
Data annotation generally refers to the labeling or generating of raw data with relevant information, which could be used for improving the efficacy of machine learning models. The process, however, is labor-intensive and costly. The emergence of LLMs presents an unprecedented opportunity to automate the complicated process of data annotation by LLM-as-a-Judge. Most of the data need to be evaluated by LLM-as-a-Judge is generated by models, or large-scale crawled data. Language models first conduct supervised fine-tuning to imitate how to align with human instructions (Wang et al., 2023a; Taori et al., 2023). After that, reinforcement learning techniques have been explored to align language models with human preferences (Ouyang et al., 2022; Ramamurthy et al., 2023). The most successful way is applying a RLHF framework (Ouyang et al., 2022) via training a reward model on human feedback and using PPO (Schulman et al., 2017) to obtain the policy model for language generation. However, in practices, the PPO training paradigm is complex in coding and hyper-parameter tuning while it needs four models that are hard for training. This motivates us to explore simpler and more straightforward methods to align language models with human preferences. This involves how to use LLM-as-a-Judge to evaluate whether different responses are aligned with human preferences. For example,  (Yuan et al., 2023b; Dong et al., 2023) use general LLM (ChatGPT) to get better alignment with human preferences. The Aplaca prompts (Taori et al., 2023) is used as sampling queries to different models generate responses. And these data was evaluated by LLM-as-a-Judge to obtain human preference scores (reward score) to train a new language model. Other works would like to use Supervised Fine-Tuning (SFT) model itself as evaluator, like generating better-aligned datasets for SFT including hindsight-modified prompts (Zhang et al., 2023a; Liu et al., 2023d) and principle-driven self-alignment (Sun et al., 2023b).
In addition, the lack of domain-specific model training data is a common phenomenon. In order to obtain annotated high-quality data, it is also very common to use LLM-as-a-Judge for the generation and evaluation of domain data. WizardMath (Luo et al., 2023) would use its Instruction Reward Model (IRM) as Evaluator, aiming to judge the quality of the evolved instructions on three aspects: i) Definition, ii) Precision, and iti) Integrity. To produce the ranking list training data of IRM, for each instruction, ChatGPT and Wizard-E are used to generate 2-4 evolved instructions respectively. Then we leverage Wizard-E to rank the quality of those 4-8 instructions.
@@
Recent research on evaluating multimodal data focuses on addressing vision-language misalignments in Multimodal Large Language Models (MLLMs), which often cause hallucinations—outputs inconsistent with visual or contextual evidence (Li et al., 2023a; Wang et al., 2023e; Cui et al., 2023). Techniques like Reinforcement Learning from Human Feedback (RLHF) and Factually Augmented RLHF have been employed to improve model alignment by incorporating structured ground-truth data and image captions, enhancing hallucination detection (Sun et al., 2023a). Benchmarks such as MLLM-as-a-Judge (Chen et al., 2024b) assess these models using tasks like scoring, pair comparison, and batch ranking, revealing limitations in alignment with human preferences. Persistent issues include biases (e.g., position, verbosity) and hallucinations, with even advanced models like GPT-4V displaying challenges. While pair comparison tasks align better with human judgment, scoring and batch ranking require significant improvements for reliable deployment. These findings emphasize the need for innovative frameworks and datasets to refine MLLM evaluation and alignment.
@@ 2.4.3. LLM-as-a-Judge for Agents
There are two ways to apply LLM-as-a-Judge for an agent. One is to evaluate the entire process of the intelligent agent (Zhuge et al., 2024), and the other is to evaluate it at a specific stage in the agent framework process (Hao et al., 2023; Shinn et al., 2023). Both approaches are briefly illustrated in Figure 10. Using LLM as the brain of agent, an agentic system (Zhuge et al., 2024) could evaluate like a human, it would reduce the need for human involvement and eliminate the trade-off between thoroughness and effort. In addition, the agent (Shinn et al., 2023) can interact with the environment through language and receive feedback on actions through LLM to make decisions for the next action.
@@ 2.4.4. LLM-as-a-Judge for Reasoning/Thinking
The role of LLM-as-a-Judge in enhancing reasoning or thinking can be understood through two frameworks: scaling training time (Gao et al., 2023c; Trung et al., 2024) and scaling test time (Snell et al., 2024). In the training phase, LLM-as-a-Judge frequently operates within reinforcement learning paradigms, where it functions as a reward model or evaluator for data or processes. This enables the creation of high-quality reasoning datasets through mechanisms such as step-by-step verification (Lightman et al., 2023), Direct Preference Optimization(DPO) (Rafailov et al., 2024), and self-refinement (Yuan et al., 2024). Recently, several LLMs trained with reinforcement learning to exhibit advanced reasoning and thinking abilities have gained attention, such as o1222https://openai.com/index/learning-to-reason-with-llms/, DeepSeek-R1333https://api-docs.deepseek.com/news/news1120,gemini-thinking [^4] https://ai.google.dev/gemini-api/docs/thinking-mode, and QVQ [^5] https://huggingface.co/Qwen/QVQ-72B-Preview. In the test-time framework, LLM-as-a-Judge is crucial for evaluating and selecting the best reasoning paths. For example, in ”Best-of-N” generation scenarios, where multiple reasoning outputs are produced, the judge determines the most accurate and coherent response. This dual role in both training and test phases demonstrates the indispensable nature of LLM-as-a-Judge in enhancing reasoning systems.
@@ 3. Improvement Strategy
When directly utilizing LLMs to conduct evaluation tasks—such as scoring, selection, pairwise comparison, or ranking—their inherent biases of LLMs like length bias, position bias, and concreteness bias(Park et al., 2024) will undermine evaluation outcomes. Mitigating these inherent biases and improving the overall evaluation performance of LLMs remains a critical challenge for applying LLMs as evaluators. In this section, we introduce three improvement strategies to boost the evaluation performance of LLM-as-a-judge: design strategy of evaluation prompts (in-context learning based), improvement strategy of LLMs’ evaluation capabilities (model-based), and optimization strategy of final evaluation results (post-processing based). As shown in Figure 12, our categorization is based on the formal definition of LLM-as-a-judge in Section 2, focusing on enhancing the evaluation effectiveness by targeting three key phases of the process: the context $\mathcal{C}$, the abilities of LLMs themselves $\mathcal{P_{LLM}}$ and the post-processing $\leftarrow$ to obtain the final results $\mathcal{E}$
@@
Evaluation Tasks, text width=6em [Few-shot prompting: FActScore (Min et al., 2023) / SALAD-Bench (Li et al., 2024b) / GPTScore (Fu et al., 2024)] [Evaluation steps decomposition: G-Eval (Liu et al., 2023b) / DHP (Wang et al., 2024f) / SocREval (He et al., 2024) / BSM (Saha et al., 2024)] [Evaluation criteria decomposition: HD-Eval (Liu et al., 2024a) / Hu and Gao et al. (Hu et al., 2024)] [Shuffling contents: Wang et al. (Wang et al., 2024c) / Auto-J (Li et al., 2023b) / JudgeLM (Zhu et al., 2023a) / PandaLM (Wang et al., 2023c)] [Conversion of evaluation tasks: Liu et al. (Liu et al., 2024b)] ] [Optimizing LLMs’
Output Forms, text width=6em [Constraining outputs in structured formats: G-Eval (Liu et al., 2023b) / DHP (Wang et al., 2024f) / LLM-EVAL (Lin and Chen, 2023)] [Providing evaluations with explanations: CLAIR (Chan et al., 2023b) / FLEUR (Lee et al., 2024b)] ] ] [Improvement Strategy
@@
Evaluation Datasets, text width=7em [PandaLM (Wang et al., 2023c) / SALAD-Bench (Li et al., 2024b) / OffsetBias (Park et al., 2024) / JudgeLM (Zhu et al., 2023a) / CritiqueLLM (Ke et al., 2024)] ] [Iterative Optimization
Based on Feedbacks, text width=7em [INSTRUCTSCORE (Xu et al., 2023b) / JADE (Zhang et al., 2023c)] ] ] [Optimization Strategy
@@
Evaluation Results, text width=7.5em [Summarize by multiple rounds: Sottana et al. (Sottana et al., 2023) / PsychoBench (Huang et al., 2023) / Auto-J (Li et al., 2023b)] [Vote by multiple LLMs: CPAD (Liu et al., 2023e) / Bai et al. (Bai et al., 2023)] ] [Direct Optimization of
LLMs’ Outputs, text width=7.5em [Score smoothing: FLEUR (Lee et al., 2024b) / G-Eval (Liu et al., 2023b) / DHP (Wang et al., 2024f)] [Self validation: TrueTeacher (Gekhman et al., 2023)] ] ] ]
@@ 3.1. Design Strategy of Evaluation Prompts
An evaluation prompt is an input to LLM evaluators, which is used to guide the LLMs to complete the required evaluation tasks. LLMs possess in-context learning ability, enabling them to learn how to perform specified tasks from relevant examples or instructions in prompts, without requiring weight updates or retraining(Brown et al., 2020). This suggests that the design strategy of evaluation prompts will significantly impact the effectiveness of LLM-as-a-judge. Therefore, optimizing the design of evaluation prompts, including better methods to help LLMs interpret the evaluation tasks and generate results, is the most direct and effective way to boost the evaluation performance of LLM-as-a-judge.
@@ 3.1.1. Optimizing LLMs’ Understanding of Evaluation Tasks
In optimization methods of prompting LLMs to better understand evaluation tasks, one of the most commonly used and effective approaches is few-shot prompting(Brown et al., 2020). By incorporating several high-quality evaluation examples into the evaluation prompts, LLM evaluators can effectively grasp the objectives, general processes, and rough evaluation criteria of evaluation tasks. Many research works employ this prompt paradigm for evaluation, such as FActScore(Min et al., 2023), SALAD-Bench(Li et al., 2024b), and GPTScore(Fu et al., 2024).
@@
In addition to providing high-quality examples for LLMs, refining the evaluation task instructions is also an effective approach to optimize LLMs’ understanding of evaluation tasks. Current methods for refining evaluation tasks mainly include the decomposition of evaluation steps and criteria: (a) Decomposition of Evaluation Steps entails breaking down the entire evaluation tasks into smaller steps, providing detailed definitions and constraints for each small step in prompts, thereby guiding LLMs comprehensively through the whole evaluation pipeline. For instance, G-Eval(Liu et al., 2023b) and DHP(Wang et al., 2024f) use Chain-of-Thought(CoT)(Wei et al., 2022) to guide LLMs. SocREval(He et al., 2024) employs the Socratic method to meticulously design each step to enhance evaluation performance. Saha et al. propose Branch-Solve-Merge(BSM)(Saha et al., 2024), which divides evaluation tasks into multiple parallel sub-tasks for separate evaluation and final merge. (b) Decomposition of Evaluation Criteria involves breaking down coarse evaluation criteria like Fluency into finer-grained sub-criteria like Grammar, Engagingness, and Readability, and then generating overall scores based on these different dimensions. HD-Eval(Liu et al., 2024a) iteratively aligns LLM evaluators with human preference via hierarchical criteria decomposition thereby addressing the potential bias in LLMs. Hu and Gao et al.(Hu et al., 2024) summarize and clearly define an explicit hierarchical classification system encompassing 11 criteria, addressing the issue of LLMs potentially confusing different evaluation standards. These refinements are specific to enable LLMs to understand the details of evaluation tasks more deeply, thereby aligning evaluation results more closely with human evaluation requirements and preferences.
@@
Furthermore, the evaluation capabilities can be optimized based on specific shortcomings of LLMs in prompts. For instance, to address specific biases like position bias which is common in pairwise evaluations, several research efforts have optimized prompts design by randomly swapping contents to be evaluated. Wang et al.(Wang et al., 2024c) analyzed and validated the impact of position bias on LLM-as-a-judge and proposed a calibration framework to mitigate this bias by swapping the contents and averaging the scores. Auto-J(Li et al., 2023b) and JudgeLM(Zhu et al., 2023a) also enhance the evaluation consistency by shuffling the texts to be evaluated. In contrast to averaging scores, PandaLM(Wang et al., 2023c) annotates the conflicting evaluation results after swapping as ”Tie” to address the position bias.